# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ 2.7.1 Radio Luminosity
One of the primary factors determining whether a pulsar is observable is its radio luminosity. Past population studies have shown that the radio luminosity of a pulsar has no dependency on $P$ or $\dot{P}$ of the pulsar itself (Szary et al., 2014). Instead, in the unit of mJy kpc2 in 1400 MHz band, it follows a log-normal distribution (Szary et al., 2014)
@@
If a pulsar crosses the “death-line”, described in § 2.5, before the observation is made, then the pulsar is not going to be observed. Additionally, another similar criterion is to examine the radio emission efficiency (Szary et al., 2014) of a pulsar, defined as the fraction of rotational energy transformed into radio emission
$\xi=\frac{L_{\textrm{r}}}{\dot{E}},$ (14)
where $L_{\textrm{r}}(\textrm{erg s}^{-1})\approx 7.4\times 10^{27}L_{\textrm{r}}$ (mJy). As the pulsar ages, its radio emission efficiency will increase. For a pulsar to be detectable in radio, $\xi<0.01$ must be satisfied.
@@ 2.7.2 PSREvolve: Radio Sensitivity
To determine if a pulsar can be observed by a given radio survey, we use the PSREvolve code (Osłowski et al., 2011; Chattopadhyay et al., 2020) to calculate the radiometer equation (Dewey et al., 1985; Lorimer & Kramer, 2004) and determine the minimum detectable flux, $S_{\textrm{min}}$, at the sky location at a given signal-to-noise ratio $(S/N_{\textrm{min}})$
@@
where $T_{\textrm{rec}}$ and $T_{\textrm{sky}}$ are the temperature of the receiver and the direction in the sky, respectively; $\Delta f$ is the bandwidth of the receiver, $G$ is the gain of the telescope, $n_{p}$ is the number of polarisations in the detector, $W_{e}$ is pulse width and $P$ is the period of the pulsar, and $\beta$ is a noise correction factor that increases the noise with digitisation errors and bandpass distortion.
Pulsar detection has been carried out by almost every major radio facility in the world, over a period of more than 50 years, using a variety of different backends. This makes it complicated to individually model each pulsar survey that has been conducted. However, one can focus on modelling some of the largest, most successful surveys (e.g., Faucher-Giguère & Kaspi, 2006; Graber et al., 2023). In this work, we choose pulsar survey parameters for Equation 15 to represent the Parkes Multibeam Pulsar Survey (Manchester et al., 2001; Lorimer et al., 2006; Sengar et al., 2023) by setting $\beta=1$ and $S/N_{\textrm{min}}\geq 10$. This survey discovered over half of the pulsars detected. Additionally, when comparing the properties of radio pulsars observed by Parkes multibeam surveys and the entire radio pulsar population, the difference in their distributions is negligible, as shown in Fig. 1. For these reasons, we assume that the radio selection is based on a “global” Parkes survey that covers the entire sky, and we use the same specifications as Parkes.
Figure 1: Cumulative distribution functions (CDFs) of $\dot{E}$ , radio flux, period and $\dot{P}$ of all radio pulsars in the catalogs (red) and pulsars observed in the Parkes Multibeam Surveys (blue).
The expression of $W_{e}$ below (Burgay et al., 2003) describes how the ISM impacts the observed pulses. Free electrons in the Galaxy broaden the intrinsic pulses $W_{i}$ (Cordes & Lazio, 2002), and the interstellar medium scatters the pulsar beam.
@@
where $W_{i}$ is the intrinsic width of the pulse, $\tau_{\textrm{samp}}$ and $\tau_{\textrm{scatt}}$ sampling and ISM scattering times, DM is the dispersion measure in the direction of the pulsar and DM0 is the diagonal dispersion measure of the survey. While the duty cycle of pulsars varies widely, here we assume $\frac{W_{i}}{P}=0.05$ for all pulsars. If the radio luminosity of a pulsar is larger than $S_{\textrm{min}}d^{2}$, then it can be detected by the chosen radio survey.
@@ 2.7.3 Radio beaming
The radio beaming fraction of a pulsar, $f_{r}$ (Tauris & Manchester, 1998), is considered to be
@@ 2.8.1 Gamma-ray Luminosity & Flux
An empirical description of the relationship between the $\gamma$-ray luminosity of a pulsar and its physical properties as a fundamental plane has been proposed (Kalapotharakos et al., 2019, 2022). In this work we estimate pulsar $\gamma$-ray luminosity using a fundamental plane in four-dimensional (4D) space that is fitted to 4FGL-DR3 pulsars (Kalapotharakos et al., 2022),
@@
where $E_{\textrm{cut}}$ is the cutoff energy of a pulsar’s $\gamma$-ray SED when modelled by a power-law with the exponential cutoff model, $B$ is the surface magnetic field strength, and $\dot{E}$ is the spin-down luminosity. Pulsar $\gamma$-ray spectral parameters, cutoff energy $E_{\textrm{cut}}$ and power-law index before cutoff $\Gamma$, are determined following Figs. 10 & 11 in Kalapotharakos et al. (2022). These two figures show correlations between $E_{\textrm{cut}}$ and $\dot{E}$  and $\Gamma$ and $\dot{E}$ respectively when the pulsar has $10^{33}$ erg/s $<$ $\dot{E}$ $<10^{38}$ erg/s. Within this $\dot{E}$ range, we assume a power-law correlation of $E_{\rm cut}$ (in GeV) with $\dot{E}$  as well as $\Gamma$ with $\dot{E}$  with these following correlations
@@
where $U(0.05)$ is a random number drawn from a uniform distribution between -0.05 and 0.05. When $\dot{E}$ $<10^{33}$ erg/s, we set
@@
where $d$ is the distance of the pulsar to the Earth and $c_{g}$ is the flux correction factor (Watters et al., 2009), which depends on the emission mechanism of the pulsar as well as geometry and viewing angle. Here we combine results from Johnston et al. (2020) and Pétri (2011), and assume that $c_{g}$ follows a sigmoid function versus $\dot{E}$ with a maximum of 0.8 (Johnston et al., 2020) and a minimum of 0.22 (Pétri, 2011), while turning over at $10^{33}$erg s-1 with an added randomisation drawn from a uniform distribution between -0.1 and 0.1.
@@ 2.8.2 Gamma-ray Sensitivity
The 12-year Fermi-LAT Source Catalog provides an all-sky map of the detection threshold for the catalog [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/12yr_catalog/detthresh_P8R3_12years_PL22.fits. By comparing the energy flux of a pulsar in the simulation to the detection threshold at the corresponding sky location, we would be able to determine if the simulated pulsar would be detectable or not.
In reality, pulsars that are already detected in radio surveys first are easier to find in $\gamma$-rays when compared to $\gamma$-ray blind searches (Abdo et al., 2013; Smith et al., 2023). Fig. 17 of Abdo et al. (2013) clearly indicates that for the detected pulsars in the 2nd Fermi-LAT pulsar catalog, pulsars that are already detected in radio in general are easier to detect. For this reason, we make the detection threshold five times lower at the location of the pulsar if the simulated pulsar is detected in radio selection; and 1.5 times higher at the location of the pulsar if the simulated pulsar is not detected in radio.
@@ 2.8.3 Gamma-ray beaming
Gamma-ray pulsars have a beaming effect similar to that of radio pulsars. In Johnston et al. (2020), a description of young and energetic (with $\dot{E}$ $>10^{35}$ erg/s) pulsars’ $\gamma$-ray beaming fraction ($f_{g}$) is in their Table. 3. We adopt this description in determining the beaming fractions of our modelled pulsars. As this work does not provide a beaming fraction for pulsars with $\dot{E}$ $<10^{35}$ erg/s, here we assume that as $\dot{E}$ decreases, the beaming fraction also decreases. The choices of beaming fraction can be described as
@@
The assumptions above are made based on observations and outer gap model of young, powerful pulsars (Johnston et al., 2020). According to Pierbattista et al. (2012), this choice can be dependent on the pulsar emission model and is consistent with what we have chosen.
@@ 3 Results & Discussions
The results from the COMPAS simulations using the setups from § 2 are presented in this section, as well as the discussion on the implications of these results. The results are compared to the pulsar catalog to determine the goodness of the different models. This will help us determine the model of pulsar initial conditions that best fits both the radio and $\gamma$-ray observations. We will first show the results of the simulation by comparing the number of different types of pulsars produced in different simulation models, then employ a statistical method to compare the distribution of pulsar physical properties to those reported in the catalogs. We further discuss the impact of the results from the selection effects of $\gamma$-ray pulsars. We also aim to provide some insights into the $\gamma$-ray emission of low $\dot{E}$ pulsars.
3.1 Star Formation Rate
In COMPAS simulations produced for this study, only stars within the initial mass range of 5 M to 150 M are modelled. Because of this setup, we need to account for the low-mass end of the initial mass function (IMF) when calculating the star formation rate. Since all the models in this study use the same parameters for binary population, they all have the same star formation rate. Using the CosmicIntegration Python package in COMPAS, we are able to account for the lower end of the IMF. By default, we use a Kroupa (2001) broken power-law IMF, with a slope of 0.3 between 0.01 M and 0.08 M ; slope of 1.3 between 0.08 M and 0.5 M ; slope of 2.3 between 0.5 M and 200 M . Each binary system generated in any given COMPAS run corresponds to 234 M of star formation. Considering we set all the binary systems to be born in the most recent 200 Myr, and each COMPAS output is sampled 10 times, the star formation rate is $\sim 1.2$ M per year with the current model selection. The observed star formation rate over the most recent history of the Milky Way is around $1.65\pm 0.19$ (Licquia & Newman, 2015) or $1.9\pm 0.4$ Myr-1 (Chomiuk & Povich, 2011). Other works, such as Evans et al. (2022), estimate the Milky Way star formation rate in the range of 0.50 - 5.93 M yr-1. The star formation rate from our COMPAS simulations is consistent with these observed and modelled values.
3.2 Results
The most straightforward and obvious way to determine if a COMPAS model is a good fit for the observation is to compare the number of pulsars produced by the model to different catalogs. We examine, for each model, the number of radio pulsars, the number of $\gamma$-ray pulsars, the number of radio-quiet $\gamma$-ray loud pulsars, and the number of radio-loud $\gamma$-ray quiet pulsars, produced in the simulation. To fully eliminate the possible randomness from the models, the observation part of post-processing is repeated 10 times, and the average and standard deviation of these realisations are recorded for each model in Table. 3.
Table 3: Results of the simulations: number of different types of pulsars. Columns left to right: COMPAS model; pulsars that are detected in radio ($N_{\rm R}$); pulsars detected in $\gamma$-ray ($N_{\rm G}$); radio-quiet $\gamma$-ray loud pulsars ($N_{\rm RQGL}$); radio-loud $\gamma$-ray quiet pulsars ($N_{\rm RLGQ}$); radio-loud $\gamma$-ray pulsars ($N_{\rm RLGL}$). The first row lists the total number of pulsars in the radio and $\gamma$-ray catalogs (Manchester et al., 2005; Smith et al., 2023). MSPs are excluded as we focus on canonical pulsars. For the simulations, the values and uncertainties are obtained from the average and standard deviation of 10 different realisations of each model.
Model NR NG NRQGL NRLGQ NRLGL
Catalogs 2715 160 90 2645 70
INIT $3414\pm 46$ $195\pm 13$ $113\pm 6$ $3331\pm 47$ $82\pm 10$
D1 $3412\pm 55$ $188\pm 7$ $103\pm 8$ $3327\pm 53$ $85\pm 7$
D2 $2533\pm 25$ $208\pm 8$ $127\pm 6$ $2451\pm 29$ $81\pm 9$
D3 $3357\pm 40$ $194\pm 6$ $114\pm 6$ $3277\pm 40$ $80\pm 7$
B1 $2803\pm 56$ $185\pm 10$ $108\pm 4$ $2726\pm 57$ $77\pm 8$
B2 $20466\pm 89$ $612\pm 18$ $199\pm 7$ $20053\pm 87$ $413\pm 14$
P1 $1953\pm 34$ $41\pm 4$ $22\pm 2$ $1935\pm 35$ $18\pm 2$
P2 $3366\pm 45$ $210\pm 10$ $112\pm 5$ $3269\pm 47$ $98\pm 8$
P3 $1993\pm 26$ $35\pm 4$ $18\pm 3$ $1976\pm 27$ $16\pm 3$
PN1 $3263\pm 47$ $138\pm 9$ $75\pm 4$ $3200\pm 45$ $62\pm 6$
PN2 $3436\pm 64$ $190\pm 9$ $109\pm 6$ $3356\pm 60$ $81\pm 7$
PN3 $3337\pm 59$ $188\pm 8$ $102\pm 7$ $3251\pm 53$ $87\pm 8$
BLN1 $3434\pm 58$ $184\pm 10$ $106\pm 7$ $3356\pm 58$ $78\pm 6$
BLN2 $12020\pm 93$ $451\pm 14$ $179\pm 12$ $11748\pm 93$ $272\pm 13$
BLN3 $1773\pm 28$ $185\pm 8$ $124\pm 7$ $1711\pm 26$ $62\pm 7$
BLN4 $7918\pm 74$ $373\pm 13$ $177\pm 7$ $7723\pm 70$ $196\pm 10$
PN1-B1 $2736\pm 27$ $146\pm 7$ $79\pm 5$ $2669\pm 26$ $67\pm 6$
BLN1-PN1 $3446\pm 59$ $152\pm 5$ $75\pm 4$ $3369\pm 59$ $77\pm 5$
BLN1-PN5 $2586\pm 34$ $52\pm 5$ $25\pm 3$ $2559\pm 33$ $27\pm 5$
BLN4-PN1 $7727\pm 79$ $262\pm 13$ $96\pm 4$ $7562\pm 80$ $166\pm 13$
We then tabulate the difference of numbers of pulsars in Table. 4. For many of the models, the predicted number of pulsars is largely different from the observed pulsars. For example, models INIT and D1 have 26% differences in radio pulsars, and 21% and 18% differences in $\gamma$-ray pulsars, respectively. An even more extreme case is B2 model, with 7.5 times of radio pulsars and 3.8 times of $\gamma$-ray pulsars predicted compared to the observed populations. Due to these large differences, we scale the predicted numbers according to the radio pulsar population. For each model, the numbers of all types of pulsars are scaled by the factor ${\rm N_{\rm R}}/2715$, with 2715 is the number of observed radio pulsars from the catalogs. These scaled numbers of $\gamma$-ray pulsar, radio-quiet and radio-loud $\gamma$-ray pulsars, and radio-loud $\gamma$-ray-quiet pulsars are then compared to those from the catalogs to calculate the percent difference. The B1 and PN1-B1 models are not scaled as the predicted numbers of radio pulsars are already close to the observed value. Scaling the number of pulsars in a model effectively changes the modelled star formation rate. We scale the default star formation rate of 1.2M /yr, as described in § 3.1, by the same scaling factor for each model, to obtain the appropriate number as listed in the last column in Table. 4. Another indicator of the goodness of the models is the ratio between different types of pulsars. Here we check the ratio between radio and $\gamma$-ray pulsars, and ratio between radio-quiet and radio-loud $\gamma$-ray pulsars for each model, and compare them to those from the catalogs. The percent differences between these predicted and observed ratios are also listed in Table. 4. Both the INIT and PN1-B1 models are able to reproduce the observed numbers and ratios of different types of pulsars to within $\approx 10\%$.
Table 4: Percentage difference (PD) between the numbers of different types of pulsars generated in COMPAS models and the radio and $\gamma$-ray catalogs. For all the models, except for B1 and PN1-B1, the comparisons are carried out after scaling the number of radio pulsars to be the same as the catalog. First five columns: percent differences of the numbers of radio, $\gamma$-ray , radio-quiet $\gamma$-ray-loud, radio-loud $\gamma$-ray-quiet, and radio-loud $\gamma$-ray-loud pulsars between each model and the catalogs. Sixth column: percent difference of the ratio between radio and $\gamma$-ray pulsars between each model and the catalogs. Seventh column: percent difference of the ratio between radio and $\gamma$-ray pulsars between each model and the catalogs. Last column: star formation rate of each model after scaling by the number of radio pulsars.
Model PDR PDG PDRQGL PDRLGQ PDRLGL PDQ1 PDQ2 SFR (M/yr)
INIT - 3.1% 0.0% 0.11% 7.14% 7.01% 1.2% 0.95
D1 - 6.25% 8.9% 3.97% 2.8% 6.95% 5.5% 0.95
D2 - 39.4% 51.1% 0.68% 24.3% 28.2% 22.6% 1.3
D3 - 1.9% 2.2% 0.2% 7.1% 1.9% 11.3% 0.97
B1 3.24% 15.6% 20% 3.1% 10% 10.72% 9.37% 1.2
B2 - 49.3% 71.1% 0.6% 21.4% 97% 62.3% 0.16
P1 - 64.4% 65.6% 1.7% 64.3% 180.5% 4.7% 1.67
P2 - 5.6% 0.0% 0.3% 12.8 5.6% 11% 0.97
P3 - 70% 72.2% 1.7% 68.6% 235.3% 12.5% 1.6
PN1 - 28% 31% 0.64% 27% 39.30% 5.47% 1.0
PN2 - 6.9% 5% 0.26% 9.3% 6.54% 4.69% 0.95
PN3 - 5% 7.8% 0.0% 1.4% 4.6% 8.6% 0.98
BLN1 - 9.4% 6.7% 0.3% 11.4% 9.9% 6.2% 0.95
BLN2 - 36.9% 55.5% 0.3% 12.8% 57% 56.2% 0.27
BLN3 - 76.9% 111.1% 0.9% 35.7% 43.5% 56.2% 1.84
BLN4 - 20% 32.2% 0.1% 3.4% 24.9% 29.7% 0.41
PN1-B1 0.7% 8.75% 12.2% 0.9% 4.3% 10.4% 7.8% 1.2
BLN1-PN1 - 25% 34.4% 0.3% 12.8% 33.6% 24.2% 0.94
BLN1-PN5 - 65.6% 71.1% 1.5% 60% 193% 28.1% 1.26
BLN4-PN1 - 42.5% 62.2% 0.03% 17.1% 73.8% 54.7% 0.42
models not scaled.
Beyond simply comparing the total number of pulsars between observations and our simulations, we can also compare the predicted and observed distributions of various physical quantities of the pulsar populations between the catalog and the simulations. We first plot the cumulative distribution functions (CDFs) for each pulsar parameter for each of the models listed in Table 1 in Figure 2, and compare them to the observed distributions. In general, some of the COMPAS model results produce distributions in most physical quantities that match the observations. But there are a few quantities that cannot be reproduced by any of the COMPAS models (e.g., $\gamma$-ray flux).

### v2

@@ 2.7.1 Radio Luminosity
One of the primary factors determining whether a pulsar is observable is its radio luminosity. Theoretical work on modelling pulsar magnetosphere has suggested a dependence of radio luminosity on spin-down luminosity $\dot{E}$ lm2004. Recent measurements made by the MeerKAT Thousand Pulsar Array have revealed evidence for such correlation, although it is a weak correlation Posselt:2023MNRAS. Some population studies have also found no significant dependence of radio luminosity on $P$, $\dot{P}$ or $\dot{E}$ of the pulsar itself Szary2014. For simplicity, we opt to assume no correlation between radio luminosity and spin-down luminosity following Szary2014. We assume that the distribution of pulsar radio luminosities of a pulsar (in units of mJy kpc2 in 1400 MHz band) follows a log-normal distribution Szary2014
@@
If a pulsar crosses the “death-line”, described in § 2.5, before the observation is made, then the pulsar is not going to be observed. As the pulsar ages, its radio emission efficiency $\xi$ as defined in Eq. 2 will increase. As per Szary2014, for a pulsar to be detectable in radio, $\xi<0.01$ must be satisfied.
@@ 2.7.2 PSREvolve: Radio Sensitivity
To determine if a pulsar can be observed by a given radio survey, we use the PSREvolve code Oslowski2021; Chattopadhyay2020 to calculate the radiometer equation dewey1985; lm2004 and determine the minimum detectable flux, $S_{\textrm{min}}$, at the sky location at a given signal-to-noise ratio $(S/N_{\textrm{min}})$
@@
where $T_{\textrm{rec}}$ and $T_{\textrm{sky}}$ are the temperature of the receiver and the direction in the sky, respectively; $\Delta f$ is the bandwidth of the receiver, $G$ is the gain of the telescope, $n_{p}$ is the number of polarisations in the detector, $t_{\rm int}$ is the integration time, $W_{e}$ is the pulse width and $P$ is the period of the pulsar, and $\beta$ is a noise correction factor that increases the noise with digitisation errors and bandpass distortion.
Pulsar detection has been carried out by almost every major radio facility in the world, over a period of more than 50 years, using a variety of different backends. This makes it complicated to individually model each pulsar survey that has been conducted. However, one can focus on modelling some of the largest, most successful surveys (Faucher-Giguere:2006ApJ; Graber:2024ApJ, e.g.,). In this work, we choose pulsar survey parameters for Equation 15 to represent the Parkes Multibeam Pulsar Survey manchester2001 and previous binary population synthesis work Chattopadhyay2021 by setting $\beta=1$, $S/N_{\textrm{min}}\geq 10$, $T_{\textrm{rec}}=24K$, $\Delta f=$ 288 MHz, $G=0.65$ K/Jy, $n_{p}=2$, and $t_{\rm int}=2100s$. The sky temperature $T_{\rm sky}$ is determined by the location of the pulsar calculated by NIGO as described in § 2.6, and is calculated by scaling the values of sky temperature at 408 MHz from the all-sky survey by Haslam:1982A&AS to the frequency of Parkes Multibeam survey of 1374 MHz, assuming the sky temperature spectral index of -2.5 Reich:1988A&A; Hobbs:2004MNRAS; Guzman:2011A&A. The pulsar period $P$ is determined by the COMPAS simulation. The Parkes Multibeam Survey discovered around half of the pulsars detected manchester2001; Lorimer:2006MNRAS; Sengar:2023MNRAS.
The expression of $W_{e}$ below 2003Natur.426..531B describes how the interstellar medium (ISM) impacts the observed pulses. Free electrons in the Galaxy broaden the intrinsic pulses $W_{i}$ 2002astro.ph..7156C, and the ISM scatters the pulsar beam.
@@
where $W_{i}$ is the intrinsic width of the pulse, $\tau_{\textrm{samp}}=250\mu s$ the sampling time of Parkes Multibeam Survey, and $\tau_{\textrm{scatt}}$ ISM scattering times based on a fit with respect to DM from Bhat:2004ApJ, DM is the dispersion measure in the direction of the pulsar and DM0 is the diagonal dispersion measure of the survey. While the duty cycle of pulsars varies significantly 1985MNRAS.213..613L, here we assume $W_{i}/{P}=0.05$ for all pulsars for simplicity. If the radio luminosity of a pulsar is larger than $S_{\textrm{min}}d^{2}$, then it can be detected by the chosen radio survey.
Additionally, only pulsars that are located in the regions observable by the Parkes Multibeam survey are selected. Their Galactic latitude (b) and longitude (l) would satisfy
$(|b|\leq 5),\&(0\leq l\leq 50\ {\rm or}\ 260\leq l<360)$
@@ 2.7.3 Radio beaming
The radio beaming fraction of a pulsar, $f_{r}$ Tauris1998, is considered to be
@@ 2.8.1 Gamma-ray Luminosity & Flux
An empirical description of the relationship between the $\gamma$-ray luminosity of a pulsar and its physical properties as a fundamental plane has been proposed Pulsar_FP2019; Pulsar_FP2022. In this work we estimate pulsar $\gamma$-ray luminosity using a fundamental plane in four-dimensional (4D) space that is fitted to pulsars in the Fourth Fermi-LAT Source Catalogue (4FGL, 4FGL;) using 12-Year survey data (4FGL-DR3, 4FGL-DR3;), which contains 190 $\gamma$-ray detected pulsars with well defined distances in the ATNF catalogue. Pulsar_FP2022 find that the $\gamma$-ray luminosity is given by,
@@
where $E_{\textrm{cut}}$ is the cutoff energy of a pulsar’s $\gamma$-ray SED when modelled by a power-law with the exponential cutoff model, $B$ is the surface magnetic field strength, and $\dot{E}$ is the spin-down luminosity. Pulsar $\gamma$-ray spectral parameters, cutoff energy $E_{\textrm{cut}}$ and power-law index before cutoff $\Gamma$, are determined following Figures 10 & 11 in Pulsar_FP2022. These two figures show correlations between $E_{\textrm{cut}}$ and $\dot{E}$  and $\Gamma$ and $\dot{E}$ respectively when the pulsar has $10^{33}$ erg/s << $\dot{E}$ $10^{38}$ erg/s. Within this $\dot{E}$ range, we assume a power-law correlation of $E_{\rm cut}$ (in GeV) with $\dot{E}$  as well as $\Gamma$ with $\dot{E}$  simply by following the trends in Figures 10 & 11 in Pulsar_FP2022 with these following correlations
@@
where $U(0.05)$ is a random number drawn from a uniform distribution between -0.05 and 0.05. These relations were estimated visually to fit the approximate trends. Outside this $\dot{E}$ range, we assume constants for $E_{\rm cut}$ and $\Gamma$. When $\dot{E}$ $10^{33}$ erg/s, we set
@@
where $d$ is the distance of the pulsar to the Earth and $c_{g}$ is the flux correction factor Watters2009, which depends on the emission mechanism of the pulsar as well as geometry and viewing angle. Here we combine results from Johnston2020 and Petri2011, and assume that $c_{g}$ follows a sigmoid function versus $\dot{E}$ with a maximum of 0.8 Johnston2020 and a minimum of 0.22 Petri2011, while turning over at $10^{33}$erg s-1 with an added randomisation drawn from a uniform distribution between -0.1 and 0.1.
@@ 2.8.2 Gamma-ray Sensitivity
The 14-year Fermi-LAT Source Catalogue provides an all-sky map of the detection threshold for the catalogue [^4] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/detthresh_P8R3_14years_PL22.fits. By comparing the energy flux of a pulsar in the simulation with the detection threshold at the corresponding sky location, we would be able to determine if the simulated pulsar would be detectable or not.
In reality, pulsars that are already detected in radio surveys first are easier to find in $\gamma$-rays when compared to $\gamma$-ray blind searches Fermi-LAT:2013svs; Fermi-LAT:2023zzt. Figure 17 of Fermi-LAT:2013svs clearly indicates that for the detected pulsars in the 2nd Fermi-LAT pulsar catalogue, pulsars that are already detected in radio in general are easier to detect. In this work, we define $\alpha$ as the factor to raise the sensitivity floor from the default 4FGL-DR4 sensitivity when the pulsar is not detected in the radio; and $\beta$ as the factor to lower the sensitivity floor when the pulsar is detected in the radio. Evaluating Figure 17 in Fermi-LAT:2013svs, we choose $\alpha=3.0$ and $\beta=1.5$.
@@ 2.8.3 Gamma-ray beaming
Gamma-ray pulsars have a beaming effect similar to that of radio pulsars. In Johnston2020, a description of young and energetic (with $\dot{E}$ $10^{35}$ erg/s) pulsars’ $\gamma$-ray beaming fraction ($f_{g}$) is in their Table 3. We adopt this description in determining the beaming fractions of our modelled pulsars. As Johnston2020 does not provide a beaming fraction for pulsars with $\dot{E}$ $10^{35}$ erg/s, here we assume that as $\dot{E}$ decreases, the beaming fraction also decreases. The choices of beaming fraction can be described as
@@
The assumptions above are made based on observations and outer gap model of young, powerful pulsars Johnston2020. According to 2012A&A...545A..42P, this choice can be dependent on the pulsar emission model and is consistent with what we have chosen.
@@ 3 Results & Discussions
The results of the COMPAS simulations using the setups from § 2 are presented in this section, as well as the discussion on the implications of these results. The results are compared to the pulsar catalogue to determine the goodness of the different models. This will help us determine the model of pulsar initial conditions that best fits both the radio and $\gamma$-ray observations. We will first employ a statistical method to compare the distribution of pulsar physical properties to those reported in the catalogues, then compare the number of different types of pulsars produced in different simulation models. These results are combined to choose a model that best describe the birth distributions of pulsar properties. We further discuss the impact of the results from the selection effects of $\gamma$-ray pulsars. We also aim to provide some insights into the $\gamma$-ray emission of low $\dot{E}$ pulsars.
3.1 Pulsar Birth Rate
In this section, we calculate the pulsar birth rate (PBR) in our COMPAS simulations. Only binaries where the primary (initially more massive) star is within the initial mass range of 5 M to 150 M are modelled using COMPAS, since lower mass stars are not expected to produce pulsars. Because of this setup, we need to account for the low-mass end of the initial mass function (IMF) when calculating the pulsar birth rate. By default, we use a Kroupa:2000iv broken power-law IMF, with a slope of 0.3 between 0.01 M and 0.08 M ; slope of 1.3 between 0.08 M and 0.5 M ; slope of 2.3 between 0.5 M and 200 M. To correct for the lower mass stars (0.08 M <M<<M< 5.0 M), we use the CosmicIntegration Python package in COMPAS to calculate a correction factor, which is the ratio of the true average mass produced per binary and the averaged mass produced per binary in COMPAS. Using the implementation of the Kroupa IMF in CosmicIntegration, we draw a sample of $10^{9}$ binaries with the mass range between 0.08 M to 150 M and calculate the average mass per binary to be $\bar{M}_{\rm true}=1.33$ M. Next we calculate the average mass per binary in COMPAS simulations. Since all the COMPAS models in this study use the same parameters for initial mass and orbit for the binaries and the same evolution physics, they all have the same pulsar birth rate. In each COMPAS simulation, the total mass simulated in binaries within the 5 to 150 M range is $M_{\rm tot,COMPAS}=8.05\times 10^{6}$ M. The average mass formed per binary in one COMPAS simulation is $\bar{M}_{\rm COMPAS}=20.3$  M. The correction factor $\kappa$ is calculated as $\kappa=\bar{M}_{\rm COMPAS}/\bar{M}_{\rm true}=15.14$.
With $N_{\mathrm{psr}}=1.2\times 10^{5}$ neutron stars produced in each simulation, the pulsar yield (the number of pulsars born per Solar mass of star formation) is calculated as,
$\zeta_{\mathrm{PSR}}=N_{\mathrm{psr}}/(M_{\rm tot,COMPAS}\times\kappa)=1.3\times 10^{-3}\,\mathrm{M}_{\odot}^{-1}.$ (18)
To translate the pulsar yield into a PBR, we have to assume a star formation history. The observed star formation rate (SFR) over the most recent history of the Milky Way is around $1.65\pm 0.19$ M yr-1 2015ApJ...806...96L, $1.9\pm 0.4$ Myr-1 2011AJ....142..197C or $3.3^{+0.7}_{-0.6}$ Myr-1 Zari:2023A&A. Other works, such as 2022ApJ...929L..18E, estimate the Milky Way star formation rate in the range of 0.50–5.93 M yr-1. The pulsar birth rate scales linearly with the assumed SFR. Here we assume the star formation rate in the recent Milky Way history is $\mathrm{SFR}=1.65$M yr-1, and the pulsar birth rate is then,
$\mathrm{PBR}=2.1\times 10^{-3}\,\mathrm{yr}^{-1}\left(\frac{\mathrm{SFR}}{1.65\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}}\right)\left(\frac{\zeta_{\mathrm{PSR}}}{1.3\times 10^{-3}\,\mathrm{M}_{\odot}^{-1}}\right)$ (19)
or 1 pulsar born every 466 years. This value is lower than most of the literature, e.g., 1/50 yr Keane:2008MNRAS, 1/46 yr to 1/58 yr Graber:2024ApJ, 1/70 yr Dirson2022, and 1/100 yr Johnston2020. The main reason for this difference could be due to some of the formation channels not included in this work, such as accretion induced collapse (AIC), stellar mergers and double white dwarf mergers. Previous COMPAS models (Stevenson:2019ApJ, e.g.,) have also found the predicted rate of core collapse supernova (CCSN) to be somewhat low compared to the observed rate. Increasing the SFR would also increase the pulsar birth rate in our work, with a detailed exploration on this effect described in § 3.5.
3.2 Results: Radio Pulsars
We start the discussion of the results of the simulations by comparing the predicted and observed distributions of various physical quantities of the pulsar populations between the catalogue and the simulations. We first discuss radio pulsars, and then discuss $\gamma$-ray pulsars in Section 3.3.
For model comparison purposes, we select Galactic radio pulsars that are not associated with any globular clusters, have positive radio flux at 1400 MHz and have an association with Parkes. Only pulsars with $P>0.03$ s and $\dot{P}>10^{-17}$ s/s are chosen, to ensure that the sample does not include pulsars that are or have been going through binary interactions/recycling. All pulsars are chosen from the ATNF Pulsar catalogue v2.5.1.
The cumulative distribution functions (CDFs) for each parameter of radio pulsars for each of the models listed in Table 1 are plotted in Figure 1555https://github.com/yuzhesong/compas_pulsar_paper.git contains all relevant plots for all models in this study.. Overall, several COMPAS models yield distributions of most physical quantities that align well with the observations. The largest variation is seen in the distributions of $P$ and $\dot{P}$.